OpenStudy (narissa):

i need help with 2 more questions

OpenStudy (studygurl14):

:)

OpenStudy (narissa):

@StudyGurl14

OpenStudy (narissa):

i will be back in 5

OpenStudy (studygurl14):

Okay, so if we do it the same way that @Jhannybean showed you... Remember that \(\large \sqrt[3]{135}=135^{\frac{1}{3}}\)

OpenStudy (studygurl14):

So, you want to split up the 135. Notice how it ends in a 5? That means that it is divisible by 5. What is 135 / 5?

OpenStudy (narissa):

27

OpenStudy (studygurl14):

Correct. So now we have... \(\large (27\times5)^{\frac{1}{3}}\)

OpenStudy (studygurl14):

Do you remember what 27 equals from the last question?

OpenStudy (studygurl14):

Like, what number when multiplied by itself three times equals 27?

OpenStudy (narissa):

4?

OpenStudy (narissa):

nooo wait let me see

OpenStudy (narissa):

its 3

OpenStudy (narissa):

so is it d?

OpenStudy (narissa):

is it d @StudyGurl14

OpenStudy (studygurl14):

Right 3. D is correct. nice job!

OpenStudy (narissa):

thank u

OpenStudy (studygurl14):

:)